(b) The following equations describe an economy being

measured in billions and I as a 5 percent interest rate.

C=0.8(1-t)Y, t=0.25, I=900-50i, Ḡ=800, L=0.25Y-62.5i,

Ḿ/Ṕ=500.

(i) What is the equation that describes the IS curve. 

2mks

(ii)What is the equation that describes the LM curve


1
Expert's answer
2021-10-04T15:25:17-0400

"i) Y=C+I+G=0.8(1-t)Y+900-50i+800\\Y=0.6Y+900-50i+800=0.8Y+1700.8-50i\\

Y-0.6Y=1700-50i\\

0.4Y=1700-50i

Y=4250-125R"

II)

"0.25Y-62.5I=500\\

0.25Y-500=62.5I\\

I=0.004Y-8"